OpenStudy (cara1382):

I have two questions I am struggling with that I have spent waaay too much time on & now it's crunch time. Any assistance would be appreciated. Form a polynomial f (x) with real coefficients having the given degree and zeros. Degree​ 4; ​ zeros: -3-2i; -4 multiplicity 2 Form a polynomial​ f(x) with real coefficients having the given degree and zeros. Degree​ 4; ​ zeros: 6​, multiplicity​ 2; 2i

OpenStudy (emilee02):

whats th answer choices

OpenStudy (sshayer):

complex zeros always occur in pairs. other zero is\[-3+2 \iota\] \[f(x)=\left\{ x-\left( -3-2 \iota \right) \right\}\left\{ x-\left( -3+2i \right) \right\}\left\{ x-\left( -4 \right) \right \}^2\]\[=\left(( x+3)+2 \iota \right)\left(( x+3)-2 \iota \right)\left( x+4 \right)^2\] first multiply first two parenthesis \[\left( a+b \right)\left( a-b \right)=a^2-b^2\] \[\left( a+b \right)^2=a^2+b^2+2ab\]

OpenStudy (mathmale):

"Form a polynomial f (x) with real coefficients having the given degree and zeros. Degree​ 4; ​ zeros: -3-2i; -4 multiplicity 2" If the zeros are -3+2i and -4, you need to expect that another zero would be -3+2i (which we call the "complex conjugate" of -3-2i. Note that there is some ambiguity here. Did you mean to say that the zeros are -3 AND 2i AND -4? Or did you mean to say that one zero is the complex -3+2i?

OpenStudy (mathmale):

Important: look up "multiplicity." How does this concept apply to the problem at hand?

OpenStudy (cara1382):

I didn't mean to say anything, the problem is written exactly how I typed it. Meaning the one zero is is the complex -3+2i

OpenStudy (cara1382):

It's not multiple choice. I think I'm messing up when I expand it. I feel like I'm doing it correctly but not getting the correct answer, so I am definitely missing something.

OpenStudy (sshayer):

\[f(x)=\left\{ \left( x+3 \right)^2-\left( 2 \iota \right)^2 \right\}(x+4)^2\] \[=\left\{ \left( x+3 \right)^2-4 \iota ^2 \right\}\left( x+4 \right)^2\] \[=\left( x^2+9+6x+4 \right)\left( x^2+8x+16 \right)\]

OpenStudy (mathmale):

"Meaning the one zero is is the complex -3+2i" Then another zero is -3-2i. Write the factors corresponding to these 2 zeros.

OpenStudy (sshayer):

simplify further

OpenStudy (mathmale):

About multiplicity: You might want to refer to the following: https://www.google.com/search?q=multiplicity&oq=multiplicity&aqs=chrome..69i57j0l5.2480j0j7&sourceid=chrome&ie=UTF-8 If the multiplicity is 2, then one root is duplicated; e. g., if one root is 2, and the multiplicity is 2, then 2 is a root twice and (x-2) is a factor twice.

OpenStudy (mathmale):

"Form a polynomial f (x) with real coefficients having the given degree and zeros. Degree​ 4; ​ zeros: -3-2i; -4 multiplicity 2" => when you've finished, you should have a polynomial whose highest x power is 4: ax^4 + bx^3 + cx^2 + d. If -3-2i is one root, then -3+2i is another root. If 2 is a root, then you have two factors: (x-2)(x-2). Write out all four factors and then multiply them together. This is what sshayer has been asking you to do.

OpenStudy (cara1382):

Will try it later, thanks everyone

OpenStudy (sshayer):



OpenStudy (sshayer):

correction last term 13*16